OpenStudy (anonymous):

Simplify: (9 – 8)2 × (24 – 6 divided by 3)

OpenStudy (anonymous):

I think the answer's 6 not sure

OpenStudy (sasogeek):

(9-8)=? multiply the answer by 2 24-6= ? divide the answer by 3 multiply the two answers...

OpenStudy (anonymous):

I got 6 as my final answer by using PEMDAS (order of operations)

OpenStudy (sasogeek):

just in case you're not sure, the answer is not 6. but this might help. 9-8=1 24-6=18 can you do the rest? (1*2) * (18/3) = what you're looking for

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks

OpenStudy (sasogeek):

is the 24-6 divided 3 the same as this \(\large 24-6 \div 3 \) ------ 1 OR this \(\huge \frac{24-6}{3} \) ----------- 2 ?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

I think I got it.............answer is 22

OpenStudy (anonymous):

i found my mistake..................I subtracted before i divided ( I'm supposed to do the reverse)

OpenStudy (sasogeek):

ok then my mistake, (9-8)*2 = ? 6/3= ? subtract your answer from 24 multiply the two answers

OpenStudy (sasogeek):

yes :)

OpenStudy (anonymous):

Thank you!